Find the cost per ounce of a mixture of 280 oz of a cologne that costs $8.50 per ounce and 560 oz of a cologne that costs $4.00 per ounce.
Answer by addingup(3677)

280+560= 840; 280/840= 0.3333; 560/840=0.6666
x= 0.3333(8.5)+0.6666(4)
x= 2.833+2.666;
x= 5.50 is the cost per ounce of your mixture